# Data Assimilation with Machine Learning Surrogate Models: A Case Study with FourCastNet

### v1

@@ 6.1 Empirical stability of 3DVar paired with FourCastNet for various observation sizes
Comparing our 3DVar analysis RMSEs and ACCs (solid lines) against our naive observation interpolation baseline (dotted lines) for each observation resolution, we notice a consistent and substantial gap in performance in favor of our 3DVar analysis. To qualitatively visualize this gap in performance, we show in Figure 2 the ground truth ERA5, interpolated 4.5 observations, and our 3DVar analysis with these 4.5 observations at the end of our assimilation window, corresponding to December 31, 2023 18:00 UTC. The interpolated observations clearly show lack of detail and are overly smooth, which is particularly noticeable in features with sharp gradients throughout the globe, such as relative humidity at 500 hPa. In contrast, our 3DVar analysis using FourCastNet and these $4.5^{\circ}$ observations show higher quality detail with an appropriate smoothness and detail given the feature. The presence of these details can be attributed to smaller-scale information encoded in the FourCastNet forward pass. To further emphasize this performance gap, we include similar visualizations comparing the ground truth ERA5 data, interpolated $4.5^{\circ}$ observations, and our 3DVar analyses with these $4.5^{\circ}$ observations for all 20 atmospheric features at the end of our year-long assimilation in Appendix D.
We note that the assimilation $5^{\circ}$ observations in Figures 1a and 1b exhibited filter divergence after assimilating about 150 days worth of observational data, despite the smoothing operation we employed in (2) for filter stability. Since a $5^{\circ}$ is a very sparse dataset, corresponding to observing only 0.25% of the latent states, and additionally given the simplifying assumptions underlying our construction of $C$ in (3), filter divergence is unsurprising. In the context of our stability theory, for this choice of $K$ that depends on $C$, the upper-bound $\lambda$ in (7) is large. We speculate that with more sophisticated assumptions on the latent state covariance $C$, analysis stability for this time horizon may be achievable.
@@
Figure 3: Visualization of various forecasting initializations for the task of $h$-step-ahead forecasting. An initialization at time $t^{\prime}$ is used to autoregressively compute forecasts up to $h$ time steps ahead using $\mathcal{F}_{\text{FCN}}$, FourCastNet. The initialization time $t^{\prime}$ varies between $1\leq t^{\prime}\leq T-h$ for all tasks (a)-(d). We consider forecasting using (a) interpolated observations, (b) true ERA5 data (unavailable in practice, serving here as a gold standard), (c) 3DVar analyses, and (d) climatology as initializations. Additionally, $t=0$ corresponds to January 1, 2023 at 00:00 UTC, and $t=T$ corresponds to December 31, 2023 at 18:00 UTC.
@@ 6.2 Forecasting accuracy given various initializations
We consider the task of $h$-step ahead forecasting given four different types of initializations. More specifically, as shown in Figure 3 we forecast with (a) interpolated $4.5^{\circ}$ observations, (b) ground truth ERA5 data (as a gold standard), (c) 3DVar analyses using these $4.5^{\circ}$ observations, and (d) climatology as initializations. Initializations (a) and (d) serve as baselines that a well-calibrated 3DVar analysis would outperform in terms of forecasting error metrics, and (b) serves as a point of comparison in order to tangibly assess the effect of the estimations in (a), (c), and (d). We report metrics in terms of standardized predictions compared to standardized ground truth ERA5 data. We utilize a climatology dataset, which corresponds to the mean value for each spatial location and feature from the years 1979 to 2015 in the ERA5 dataset.
@@
In producing Figure 4, we encountered two out of 1432 time points where our 3DVar analyses with $4.5^{\circ}$ observations, when used as initial conditions for FourCastNet, resulted in degraded forecasts after a roughly 2 day forecasting horizon. These degraded forecasts are characterized by large, physically unrealistic predicted values originating from diverging forecasts originating at a particular location on the globe. The two time points corresponded to an extreme event, specifically Typhoon Khanun, and the forecasting errors autoregressively accumulated in the region of this typhoon. Such catastrophic forecast errors were not seen in the forecasts using climatology, interpolated $4.5^{\circ}$ observations, or ground truth ERA5 data. Because we did not see the same catastrophic forecasting divergence in our 3DVar analyses with $2^{\circ}$ and $2.5^{\circ}$ observation resolutions, we hypothesize that our $4.5^{\circ}$ 3DVar analyses at these two time points do not have enough data near the typhoon to adequately estimate a physically realistic initial condition given the construction of our 3DVar algorithm, leading to downstream forecasting divergence.
@@ 6.3 Extreme event: Typhoon Mawar, 2023
We evaluate forecasts that use three different initializations: (1) our 3DVar analysis with $4.5^{\circ}$ observations, (2) interpolated $4.5^{\circ}$ observations, and, as an idealized comparison, (3) ground truth ERA5 data. Comparing our 3DVar-initialized forecasts against interpolated-observation-initialized forecasts allows us to assess the gain in performance as a result of our data assimilation framework, and comparing our 3DVar initialized forecasts against ground truth ERA5 initialized forecasts allows us to assess the performance gap in how well our 3DVar forecasts perform compared to how well we could hope to perform in an ideal scenario, where we have access to the full ground truth initial condition.
Forecasting the eye of Typhoon Mawar from May 23, 2023 to May 30, 2023.
@@
Figure 5: Visualization of the 7 day forecast of the estimated eye of Typhoon Mawar initialized on May 23, 2023 using three different initial conditions: ground truth ERA5 data as a gold standard (left), our 3DVar analysis using $4.5^{\circ}$ noisy observations (middle), and interpolated $4.5^{\circ}$ noisy observations (right). Each standardized initialization is perturbed by $\mathcal{N}(0,0.3I_{d_{x}})$ noise to create a 50 member ensemble. These initial ensemble members were then independently propagated forward in time using FourCastNet without any additional data to correct these forecasts.
@@
Our first typhoon forecasting assessment focuses on assessing the predicted location of the eye of the hurricane, which we characterize by the location of the minimum mean sea level pressure.
@@
For each of our initialization types, we add $\mathcal{N}(0,0.3I_{d_{x}})$ noise to each standardized initial estimate and create an ensemble of size 50 with these perturbations. Figure 5 visualizes the ensemble of predicted typhoon trajectories for each of our three forecasting initializations. In these plots, we include the ground truth trajectory, labeled as “Ground Truth ERA5,” to evaluate whether this ground truth trajectory is included in the ensemble spread for each initialization type.
While on the tail of the distributions, the ground truth typhoon trajectory is included in the ensemble spread for almost the full trajectory of the typhoon for both the idealized setting (initializing with noisy ground truth ERA5 data) and our setting (initializing with noisy 3DVar analyses using $4.5^{\circ}$ observations). On the other hand, initializing with noisy interpolated $4.5^{\circ}$ observations appears to forecast an ensemble spread that does to capture the ground truth trajectory as well as our 3DVar analysis initialization or the ground truth ERA5 initialization.
@@
(b) Predicted maximum wind speed at 10 meters above the surface across different initial conditions compared to the actual maximum 10 meter wind speed.
@@
Figure 6: Plots of the forecasted minimum mean sea level pressure and maximum wind speed 10 meters above the surface, initialized with ground truth ERA5 data, interpolated $4.5^{\circ}$ observations, and our 3DVar analysis using $4.5^{\circ}$ observations on May 23, 2023. The ground truth is additionally plotted to compare these forecasts. For each of our three initial conditions, we create an ensemble of size 50 by adding $\mathcal{N}(0,0.3I_{d_{x}})$ distributed noise to the standardized initializations. The forecasts were then converted back to its original scale to produce these plots. The shaded regions in both plots correspond to the 0.05 and the 0.95 quantiles of the ensemble predictions.
@@
Forecasting the maximum wind speed 10 meters above the surface and minimum mean sea level pressure.
@@
Aside from tracking the location of the typhoon, two other important ways to quantitatively characterize the typhoon include assessing the predicted minimum mean sea level pressure and maximum wind speed at 10 meters above the surface. These two features determine the categorization of the typhoon, so in order to accurately predict the intensity of the storm, predictions need to be especially accurate regarding these two features. Because wind speed is not a feature native to ERA5 data, we derived the wind speed 10 above the surface using the formula $\sqrt{\text{U10}^{2}+\text{V10}^{2}}$, where U10 and V10 correspond to the U- and V- component wind speed at 10m above the surface, respectively. Figure 6 plots forecasts for these two features across the different initializations we consider, initializing on May 23, 2023. These plots include the ground truth ERA5 minimum mean sea level pressure and maximum wind speed 10m above the surface across our forecasting horizon as a visual comparison.
@@
We note that generally, all three of our initializations under-predict the true intensity of the storm; the predicted minimum mean sea level pressure is larger than ground truth across our forecasting horizon for all three initializations, and the predicted wind speed at 10m above the surface is lower than the ground truth across our forecasting horizon for all three initializations, both of which correspond to predicting a less intense typhoon. However, especially apparent in short forecasts, our 3DVar analysis produces predicted minimum mean sea level pressure and maximum 10m wind speed closer to the “gold standard” forecasts (i.e., initializing with ground truth ERA5 data) and to the actual ground truth compared to the interpolated $4.5^{\circ}$ observation forecasts. The smoothed out features in the interpolated $4.5^{\circ}$ observations likely further contribute to under-predicting this extreme event. As the forecasting horizon increases, the advantage of our 3DVar analysis initialization appears to lessen after an approximate 2 day forecast horizon, as shown by the vast overlap in the ensemble quantiles.
@@
(a) Ground truth ERA5 (left) and 48 hour forecasted (right-most three) wind speed at 10 meters above the surface on May 25, 2023.
@@
(b) Ground truth ERA5 (left) and 48 hour forecasted (right-most three) mean sea level pressure on May 25, 2023.
@@
Figure 7: Visualization of the ground truth ERA5 wind speed at 10m above the surface and mean sea level pressure compared to 48 hour forecasts given three initial conditions: noisy ERA5 data, noisy interpolated $4.5^{\circ}$ observations, and a noisy 3DVar analysis ($4.5^{\circ}$ observations).
48 hour forecast of Typhoon Mawar, initialized on May 23, 2023.
@@
To qualitatively visualize the difference in predictions for our three initial conditions in terms of the wind speed and the sea level pressure, we provide plots of one ensemble member’s 48 hour forecasts in Figures 7a and 7b, corresponding to May 25, 2023, using our three initializations. We include the ground truth for May 25, 2023 in these figures as a visual comparison. Consistent with the conclusion drawn from Figure 6, the 48 hour forecast for interpolated $4.5^{\circ}$ observations shows a less extreme forecast in terms of wind speed and mean sea level pressure compared to the ground truth. The forecasts from initializing with the ground truth and a 3DVar analysis ($4.5^{\circ}$ observations) show a qualitatively better prediction of the eye of the hurricane, more closely matching its intensity. For a more thorough assessment across different ensemble members, we include a comparison of three other ensemble forecasts in Appendix E.

### v2

@@ 6.1 Empirical stability of 3DVar paired with FourCastNet for various observation sizes
Comparing our 3DVar analysis RMSEs and ACCs (solid lines) against our naive observation interpolation baseline (dotted lines) for each observation resolution, we notice a consistent and substantial gap in performance in favor of our 3DVar analysis. To qualitatively visualize this gap in performance, we show in Figure 2 the ground truth ERA5, interpolated 4.5 observations, and our 3DVar analysis with these 4.5 observations at the end of our assimilation window, corresponding to December 31, 2023 18:00 UTC. The interpolated observations clearly show lack of detail and are overly smooth, which is particularly noticeable in features with sharp gradients throughout the globe, such as relative humidity at 500 hPa. In contrast, our 3DVar analysis using FourCastNet and these $4.5^{\circ}$ observations show higher quality detail with an appropriate smoothness and detail given the feature. The presence of these details can be attributed to smaller-scale information encoded in the FourCastNet forward pass. To further emphasize this performance gap, we include similar visualizations comparing the ground truth ERA5 data, interpolated $4.5^{\circ}$ observations, and our 3DVar analyses with these $4.5^{\circ}$ observations for all 20 atmospheric features at the end of our year-long assimilation in Appendix F.
We note that the assimilation $5^{\circ}$ observations in Figures 1a and 1b exhibited filter divergence after assimilating about 150 days worth of observational data, despite the smoothing operation we employed in (2) for filter stability. We visualize in Figure 10 in Appendix B the 3DVar estimate of wind speed compared to ERA5 wind speed soon after the assimilation began to exhibit instability, corresponding to May 29, 2023. The instability originates near the eye of Typhoon Mawar, and FourCastNet predicts increasingly larger wind speeds that are not adequately corrected by the sparse 5 observations in 3DVar. Since a $5^{\circ}$ is a very sparse dataset, corresponding to observing only 0.25% of the states, and additionally given the simplifying assumptions underlying our construction of $C$ in (3), filter divergence is unsurprising in this extreme case. In the context of our stability theory, for this choice of $K$ that depends on $C$, the upper-bound $\lambda$ in (7) is large. We speculate that with more sophisticated assumptions on the background covariance $C$ paired with more localized observations, analysis stability for this time horizon may be achievable.
@@
Figure 3: Visualization of various forecasting initializations for the task of $h$-step-ahead forecasting. An initialization at time $t^{\prime}$ is used to autoregressively compute forecasts up to $h$ time steps ahead using $\mathcal{F}_{\text{FCN}}$, FourCastNet. The initialization time $t^{\prime}$ varies between $1\leq t^{\prime}\leq T-h$ for all tasks (a)-(d). We consider forecasting using (a) interpolated observations, (b) true ERA5 data (unavailable in practice, serving here as an idealized setting), (c) 3DVar analyses, and (d) climatology as initializations. Additionally, $t=0$ corresponds to January 1, 2023 at 00:00 UTC, and $t=T$ corresponds to December 31, 2023 at 18:00 UTC.
@@ 6.2 Forecasting accuracy given various initializations
We consider the task of $h$-step ahead forecasting given four different types of initializations. More specifically, as shown in Figure 3 we forecast with (a) interpolated $4.5^{\circ}$ observations, (b) ground truth ERA5 data (as an ideal setting), (c) 3DVar analyses using these $4.5^{\circ}$ observations, and (d) climatology as initializations. Initializations (a) and (d) serve as baselines that a well-calibrated 3DVar analysis would outperform in terms of forecasting error metrics, and (b) serves as a point of comparison in order to tangibly assess the effect of the estimations in (a), (c), and (d). We report metrics in terms of standardized predictions compared to standardized ground truth ERA5 data. We utilize a climatology dataset, which corresponds to the mean value for each spatial location and feature from the years 1979 to 2015 in the ERA5 dataset.
@@
In producing Figure 4, we encountered two out of 1432 time points where our 3DVar analyses with $4.5^{\circ}$ observations, when used as initial conditions for FourCastNet, resulted in degraded forecasts after a roughly 2 day forecasting horizon. These degraded forecasts are characterized by large, physically unrealistic predicted values originating at a particular location on the globe. The two time points corresponded to an extreme event, specifically Typhoon Khanun, and the forecasting errors autoregressively accumulated in the region of this typhoon. A visualization of the 10m wind speed field across different forecasting horizons that show this divergence is shown in Figure 11 in Appendix B. Such catastrophic forecast errors were not seen in the forecasts using climatology, interpolated $4.5^{\circ}$ observations, or ground truth ERA5 data. Because we did not see the same catastrophic forecasting divergence in our 3DVar analyses with $2^{\circ}$ and $2.5^{\circ}$ observation resolutions, one hypothesis is that our $4.5^{\circ}$ 3DVar analyses at these two time points do not have enough data near the typhoon to adequately estimate a physically realistic initial condition given the construction of our 3DVar algorithm, leading to downstream forecasting divergence. However, we also note that some slight perturbation to the locations of assimilated observations leads to stable forecasts for these same two time points. This result suggests that small changes to the assimilated dataset, especially in a highly sparse regime, can lead to large differences in the analyses, and therefore varied forecasts when these analyses are used as initial conditions. Due to the complexity of this system and the black-box nature of machine learning surrogates, a definitive explanation for this behavior is unclear at present.
@@ 6.3 Extreme event: Typhoon Mawar, 2023
We evaluate FourCastNet’s predictions using three different initializations: (1) our 3DVar analysis with $4.5^{\circ}$ observations, (2) interpolated $4.5^{\circ}$ observations, and, as an idealized comparison, (3) ERA5 reanalysis data. Comparing our 3DVar-initialized forecasts against interpolated-observation-initialized forecasts allows us to assess the gain in performance as a result of our data assimilation framework, and comparing our 3DVar initialized forecasts against ERA5 initialized forecasts allows us to assess the performance gap in how well our 3DVar forecasts perform compared to how well we could hope to perform in an ideal scenario where we have access to a high fidelity initial condition. For additional comparison, we include forecasts from (1) the ECMWF’s IFS-HRES, which provide high resolution predictions from the IFS numerical weather model, and (2) IBTrACS observational data from the Joint Typhoon Warning Center (JTWC).
Forecasting the eye of Typhoon Mawar from May 23, 2023 00:00 UTC to May 30, 2023 12:00 UTC.
@@
Figure 5: Visualization of FourCastNet’s 7 day forecast of the estimated eye of Typhoon Mawar initialized on May 23, 2023 00:00 UTC using three different initial conditions: ground truth ERA5 data as an ideal setting (left), our 3DVar analysis using $4.5^{\circ}$ noisy observations (middle), and interpolated $4.5^{\circ}$ noisy observations (right). Each standardized initialization is perturbed by $\mathcal{N}(0,0.3I_{d_{x}})$ noise to create a 50 member ensemble. These initial ensemble members were then independently propagated forward in time using FourCastNet without any additional data to correct these forecasts. For comparison, we include the eye of the typhoon based on ERA5, a single IFS-HRES forecast, and IBTrACS observational data in each plot. The skill of the ensemble in predicting the typhoon’s trajectory based on IBTrACS data using the CRPS metric is listed at the top of each image showing the forecast ensembles.
@@
Our first typhoon forecasting assessment focuses on the predicted location of the eye of the hurricane, which we characterize by the location of the minimum mean sea level pressure.
@@
For each of our initialization types, we add $\mathcal{N}(0,0.3I_{d_{x}})$ noise to each standardized initial estimate and create an ensemble of size 50 with these perturbations. Figure 5 visualizes the ensemble of predicted typhoon trajectories for each of our three forecasting initializations. In these plots, we include two trajectories, “ERA5 Reanalysis” and “IBTrACS Data" to evaluate whether these trajectories are included in the ensemble spread for each initialization type. We note that this figure shows forecasts for initializing at only one time point, so the relative performance of initializing with ERA5, our 3DVar analysis, and interpolated observations may vary with different starting time points.
Based on Figure 5, the ensemble spread of the estimated typhoon trajectories generally contain the eye of the typhoon based on the ERA5 reanalysis and IBTrACS data for forecasts using 3DVar analysis (4.5 obs.) and interpolated 4.5 observations as initial conditions. We note that the 3DVar analysis (4.5 obs.) produces a narrower ensemble spread that appears to be closely aligned with these trajectories. However, the FourCastNet predictions initialized with ERA5 reanalysis data appear to be better calibrated with the early-time location of the typhoon’s eye; the early-stage forecasts appear to have a slight westward bias for both the 3DVar analysis and interpolated observations initial conditions. The computed continuous ranked probability scores (CRPS) for each ensemble of predictions in Figure 5 suggest that the forecasts initialized with the 3DVar analysis (4.5 obs.) and interpolated 4.5 observations are equally performing for predicting the IBTrACS observed typhoon trajectory, and both outperform initializing with ERA5 reanalysis.
All forecasts created using FourCastNet in our visualization share one trait in common: the predictions evolve the typhoon across space at a much faster rate compared to the ground truth. By comparison, the IFS-HRES prediction, despite showing some minor bias throughout the typhoon’s trajectory, has a well calibrated speed at which the typhoon moves across the space. As is similarly the case in the ERA5 trajectory, the IFS-HRES forecasts shows a slower initial-time movement, followed by a more rapid north-western movement, then again a slower pace as it dissipates.
@@
(b) Predicted maximum wind speed at 10m above the surface across different initial conditions compared to the actual maximum 10 meter wind speed.
@@
Figure 6: Plots of the forecasted minimum mean sea level pressure and maximum wind speed 10m above the surface for Typhoon Mawar, initialized with ERA5 data, interpolated $4.5^{\circ}$ observations, and our 3DVar analysis using $4.5^{\circ}$ observations initialized on May 23, 2023. The ERA5 values, IFS-HRES forecasts, and IBTrACS observations are additionally plotted to compare these forecasts. For each of our three initial conditions, we create an ensemble of size 50 by adding $\mathcal{N}(0,0.3I_{d_{x}})$ distributed noise to the standardized initializations. The forecasts were then converted back to its original scale to produce these plots. The shaded regions in both plots correspond to the 0.05 and the 0.95 quantiles of the ensemble predictions.
@@
Forecasting the maximum wind speed 10m above the surface and minimum mean sea level pressure.
@@
Aside from tracking the location of the typhoon, two other important ways to quantitatively characterize the typhoon include assessing the predicted minimum mean sea level pressure and the maximum wind speed at 10m above the surface. These two features determine the categorization of the typhoon, so in order to accurately predict the intensity of the storm, predictions need to be especially accurate regarding these two features. Because wind speed is not a feature native to ERA5 data, we derived the wind speed 10 above the surface using the formula $\sqrt{\text{U10}^{2}+\text{V10}^{2}}$, where U10 and V10 correspond to the U- and V- component wind speed at 10m above the surface, respectively. Figure 6 plots forecasts for these two features across the different initializations we consider, initializing on May 23, 2023 at 00:00 UTC. These plots include the ERA5 reanalysis, IFS-HRES forecast, and IBTrACS observational minimum mean sea level pressure and 10m maximum wind speed across our forecasting horizon as a visual comparison. The IBTrACS maximum wind speeds and minimum pressure values correspond to the most extreme 1-minute sustained values observed.
@@
We note that all three of our initializations, including the IFS-HRES forecast, under-predict the observed intensity of the storm given by the IBTrACS data; the predicted minimum mean sea level pressure is larger than the observed values across our forecasting horizon for all forecasts, and the predicted maximum wind speed at 10m above the surface is lower than the observed value for almost all time points across our forecasting horizon for all three initializations. Both of these results correspond to predicting a less intense typhoon. However, especially apparent in short-time forecasts, our 3DVar analysis produces predicted minimum mean sea level pressure and maximum 10m wind speed closer to the idealized forecasts (i.e., initializing with ground truth ERA5 data) compared to the interpolated $4.5^{\circ}$ observation forecasts. The smoothed out features in the interpolated $4.5^{\circ}$ observations likely further contribute to under-predicting this extreme event. As the forecasting horizon increases, the advantage of our 3DVar analysis initialization appears to lessen after an approximate 2 day forecast horizon, as shown by the vast overlap in the ensemble quantiles.
A notable point in Figure 6 is that the ERA5 reanalysis provides significantly less extreme values for the maximum 10m wind speed and minimum mean sea level pressure compared to the IBTrACS observational data, suggesting that the ERA5 may not capture the level of extremity observed for extreme storms.
@@
(a) Ground truth ERA5 (left) and 48 hour forecasted (right-most four) wind speed at 10m above the surface on May 25, 2023.
@@
(b) Ground truth ERA5 (left) and 48 hour forecasted (right-most four) mean sea level pressure on May 25, 2023.
@@
Figure 7: Visualization of the ground truth ERA5 wind speed at 10m above the surface and mean sea level pressure compared to 48 hour forecasts from IFS-HRES and FourCastNet’s forecasts given three initial conditions: noisy ERA5 data, noisy interpolated $4.5^{\circ}$ observations, and a noisy 3DVar analysis ($4.5^{\circ}$ observations).
48 hour forecast of Typhoon Mawar, initialized on May 23, 2023 00:00 UTC.
@@
To qualitatively visualize the difference in predictions for our three initial conditions in terms of the wind speed and the sea level pressure, we provide plots of one ensemble member’s 48 hour forecasts in Figures 7a and 7b, corresponding to May 25, 2023 at 00:00 UTC, using our three initializations. We include the ERA5 reanalysis and the IFS-HRES forecast for May 25, 2023 in these figures as a visual comparison. Consistent with the conclusion drawn from Figure 6, the 48 hour forecast for interpolated $4.5^{\circ}$ observations shows a less extreme forecast in terms of wind speed and mean sea level pressure compared to the ERA5 reanalysis. The forecasts from initializing with the ERA5 reanalysis and a 3DVar analysis ($4.5^{\circ}$ observations) show a qualitatively better prediction of the eye of the hurricane, more closely matching its intensity. For a more thorough assessment across different ensemble members, we include a comparison of three other ensemble forecasts in Appendix G. Lastly, we note that Figures 7a and 7b visually reiterate the insights that the FourCastNet predictions under-predicted the intensity of the typhoon, while the IFS-HRES forecast is closer to the extremity observed in the IBTrACS observations, though the forecast intensity still falls short of the observed intensity.